# Aesthetic Alignment Risks Assimilation: How Image Generation and Reward Models Reinforce Beauty Bias and Ideological “Censorship”

### v1

@@ Abstract
Over-aligning image generation models to a generalized aesthetic preference conflicts with user intent, particularly when “anti-aesthetic” outputs are requested for artistic or critical purposes. This adherence prioritizes developer-centered values, compromising user autonomy and aesthetic pluralism. We test this bias by constructing a wide-spectrum aesthetics dataset and evaluating state-of-the-art generation and reward models. We find that aesthetic-aligned generation models frequently default to conventionally beautiful outputs, failing to respect instructions for low-quality or negative imagery. Crucially, reward models penalize anti-aesthetic images even when they perfectly match the explicit user prompt. We confirm this systemic bias through image-to-image editing and evaluation against real abstract artworks.
1 Introduction and Related Works
@@
Following developments in Large Language Models (LLMs), many image generation models have been fine-tuned with human feedback to better align with human expectations, which is usually referred to as alignment. Alignment has two primary focuses: instruction following and general preference (aesthetics). A frequently overlooked issue is the potential conflict between these focuses: what should a model prioritize when a user request contradicts general preference? Most pipelines for general preference assume a single, universal human standard of aesthetics and quality that serves everyone’s needs, and aligning to such a preference is often treated as beneficial for safety and user experience. This is usually done by using a reward model, a model used to judge the quality of the image, as a signal to perform reinforcement learning on the generative model. This assumption appears in several reinforcement learning papers ([li_playground_2024, kim_preference_2024, liu_flow-grpo_2025]) and reward model papers ([xu_imagereward_2023, wu_human_2023, ma_hpsv3_2025, xu_visionreward_2025, kirstain_pick--pic_2023, zhang_learning_2024, wu_human_2023_1]). We agree that a mean or mode (mainstream) of general human preference exists within a population or subpopulation, merely in a statistical sense. However, we argue that strict alignment to that preference is problematic. Imposing a universal preference that overrides user instructions may undermine user autonomy and expressive agency, raising concerns about developer-centered value imposition and limiting aesthetic pluralism. What the image generation and reward models are aligned to is an imaginary, abstract person modeled by the mean preference of all Homo sapiens, not the concrete individuals of each user.
Figure 1: Luxe, Calme et Volupté, by Henri Matisse (19041904). As the origin of Fauvism, this work achieved significant artistic value by intentionally breaking the prevailing aesthetic norms of its time through anti-naturalistic color. However, despite explicit instructions for Fauvism, this image still received low HPSv3 score of 1.73 [ma_hpsv3_2025] because it does not align with mainstream notions of “good” aesthetics. In contrast, typical high-aesthetic images reach scores around 15.
@@ 1.1 The Role of Wide-Spectrum Aesthetics
In this work, “wide-spectrum aesthetics” (or anti-aesthetics) refers to images intentionally generated with low-quality attributes (e.g., blur, noise, distortion) or usually avoided patterns (e.g., unrealism, clashing color, hieratic scale) as per user instructions for experimental, critical, or technical use cases. These images might be intentionally awkward, conflicting, or unnatural. It does not mean spontaneous error without user instruction to do so or truly unsafe output (note that an image is not unsafe if it advocates against war by showing the horrors of war). The concept of aesthetics has never been clearly defined. Many artistic movements once considered unattractive, such as Fauvism (see Figure 1), Expressionism, and Abstract art, were later recognized for their artistic values. Beyond experimental art that challenges conventional notions of beauty, intentionally “ugly” art plays a crucial role in satire and social critique. As Adorno noted, “Rather, in the ugly, art must denounce the world that creates and reproduces the ugly in its own image” [sartwell_beauty_2024, adorno_aesthetic_1984]. By exposing or amplifying the flaws, injustices, and distortions of reality, such art provokes reflection and calls for change. Dadaism [noauthor_dada_nodate], which emerged during World War I, exemplifies this approach by using deliberate ugliness to confront the absurdity and horror of war.
1.2 Concerns with AI Preference Alignment
@@
Previous work has argued that a developer-set preference in LLMs for health-related queries is “unethical and dangerous” [guo_position_2025], noting that developers may prioritize legal and reputational concerns over users’ actual well-being. Other argumentative papers caution that “human value alignment” can be risky due to developer control and interests, harm to value pluralism, bias in the values being aligned to, and the possibility that human values are not inherently good [sutrop_challenges_2020, arzberger_nothing_2024, turchin_ai_2019]. Previous research has found that LLMs could have ideological bias [rozado_measuring_2025, faulborn_only_2025, buyl_large_2025, rettenberger_assessing_2025] and it could depend on their developers [buyl_large_2025], size [rettenberger_assessing_2025], or alignment process [faulborn_only_2025]. Additional details about problems with human value alignment are provided in the related work section of [guo_position_2025].
@@
1.3 Previous Discussion on Alignment of Image Generation Models
In image generation research, concerns about generalized aesthetic bias and lack of preference diversity have been raised in several studies, but not systematically argued and studied. The Value Sign Flip (VSF) pilot study [guo_vsf_2025] explored negative prompting to induce non-mainstream outputs but did not extend its findings to large-scale generative or reward models. They also did not provide a complete argument as to why over-alignment is harmful. LAPIS [maerten_lapis_2025] and HPSv3 [ma_hpsv3_2025] measured both mean and variance of human preference, yet HPSv3 continued to model general preferences rather than individual variation. Jin et al. [jin_compose_2025] proposed user-specific adapters emphasizing personalized alignment, but did not include intentionally technical degraded outputs or usually avoided patterns and did not conduct large-scale experiments on generative and reward models. The Flux Krea team [noauthor_releasing_2025] identified systematic biases in popular aesthetic reward models, arguing that averaging human values yields unsatisfactory compromises into a “no-body’s happy here” zone. HPSv3 [ma_hpsv3_2025] imposed real-world and expert-rating constraints that limit creative deviation and stylistic diversity. VisionReward [xu_visionreward_2025] decomposed human preference into interpretable sub-scores but overemphasized traits like brightness, positivity, and prominence, potentially penalizing valid low-saturation, abstract, or emotionally negative imagery, thus misaligning reward-driven models with user intent. More details are in Appendix.
1.4 Toxic Positivity
Many reward models promote strong positive emotions as an aesthetics aspect, which risk suppressing negative ones, reinforcing the false dichotomy that positive emotions are good and negative ones are bad. This view is harmful because negative emotions are integral to authentic human expression and growth. Overemphasizing positivity, as seen in social media, produces sanitized representations that obscure emotional complexity and neglect the constructive role of negative emotions. When AI-generated images consistently display happiness, they can create unrealistic emotional norms, fostering toxic positivity and discouraging healthy emotional engagement. Negative emotions also serve vital functions, such as alerting us to moral or physical danger and fostering empathy. More details and sources are in Appendix.
@@ 1.5 Previous Alignment Benchmarks
Benchmarks mirror alignment goals and generally fall into two categories: (complex) prompt following and general aesthetics. TIIF-Bench [wei_tiif-bench_2025], UniGenBench [wang_pref-grpo_2025], and GenEval [ghosh_geneval_2023] test models on complex prompt following, including spatial relationships, counting, and attributes. T2I-ReasonBench [sun_t2i-reasonbench_2025] evaluates reasoning capabilities such as idiom interpretation and real-world understanding. On the aesthetics side, many reward models report scores assigned by their own evaluators, such as ImageReward [xu_imagereward_2023], HPSv2 [wu_human_2023], and HPSv3 [ma_hpsv3_2025]. These evaluators also consider prompt following, but it remains unclear how they weigh each factor when general preference and the prompt conflict. There are also some benchmarks targeting biases in image generation models; however, they mainly focus on demographic bias and fairness and not aesthetics aspects [seshadri_bias_2023, wan_survey_2024].
1.6 Five Layers of Concerns
We proposed that the risks associated with imposing this universal preference can be analyzed through five distinct layers of ethical and practical debate:
@@ 1.6 Five Layers of Concerns
Developer’s or Users’ Preference The question here is whether the alignment a developer implements truly promotes genuine human-centered values (for the user’s good), or if it primarily serves the developer’s own benefit, such as mitigating reputation, legal, or marketing risks (developer-centered value) [guo_position_2025]. We argue that this pre-emptive exclusion of non-mainstream outputs, driven by developer values, constitutes pre-emptive governance [lazar_governing_2025]. This modality of power, exercised through algorithmic design, challenges the political-philosophical notion of authority and undermines relational equality by unilaterally deciding the terms of creative possibility. For instance, when an AI avoids generating critical art, is it protecting the company or the user? This practice effectively eliminates the user’s resistibility—a critical democratic safeguard—by designing away the option to dissent from the system’s imposed aesthetic norm.
@@
Inherit Bias Even if the developer is not self-interested, their view of human preference may still reflect their own values and background, leading to a well-intentioned yet narrow definition of “good” that overlooks aesthetic diversity. Research shows that AI models tend to encode and amplify dominant beauty standards, frequently biasing generated images towards Western features and excluding non-normative representations [vargas2025visual]. This effect is often seen through the active removal of features thought of as “undesirable” or “ugly”, which further propagates the beauty myth in generative outputs[dinkar_erasing_2025]. This phenomenon comes from training data showing the tastes of specific demographics, which solidifies a limited cultural capital and results in the homogenization of aesthetic output [vianna2025aesthetic]. Consequently, the quantification of beauty by AI, while it looks “fair”, ultimately risks cultural differences and weakens the concept of diversity [chen2024study]. However, existing work mainly focuses on the bias from demographic and culture bias. We argue here that these biases also include general preferences like lighting, color, styles, unrealism, clashing color, hieratic scale, etc.
People’s preferences or a person’s preference The third question asks, even if the generalized “good” is truly beneficial for most users (a human-centered value), can this generalized standard rightly replace the user’s own specific intent (a person-centered value)? When users explicitly request “low-quality” images and the model automatically cleans them up, accuracy to the user’s intent is also lost. It is similar to a pineapple sorter that discards anything not matching the general preferred flavor, even when the user deliberately orders something sour or an unusual pineapple flavor.
The problem of excessive single value Overzealous adherence to a single preference or subset of preferences could result in unpredictable harm. An image generation AI might tend to overly represent bright colors, and chasing universal preference might actually fail outside of generalized preference because of the extremely bright colors. This is also related to the virtue theory from ancient Greece, where moderation of any characteristic is important [guo_position_2025].
@@
The problem of sanitized reality When an image generator produces outputs that are polished, flawless, and universally beautiful, does it still reflect reality or the user’s intent? If every image resembles an idealized Instagram wonderland, it risks becoming a fantasy rather than a mirror of truth, echoing the artificial harmony of Brave New World.
1.7 Our Contribution
Our contributions are: (1) We argued that while a general mean or mode of human aesthetic preference exists, over-aligning generative models to it hurts user control, especially when users explicitly request non-mainstream or intentionally “ugly” outputs. This differs from prior work on cultural or demographic bias in beauty standards because we focus on purposely decreasing technical and aesthetic quality to achieve wide-spectrum aesthetic effects for experimental, critical, or technical uses. (2) To test this, we built a dataset of wide-spectrum aesthetics prompts and evaluated whether current image generators can reliably produce such content, comparing base models with their aligned variants (aesthetics- or instruction-tuned) to see whether alignment harms the ability to generate wide-spectrum aesthetics images. (3) We also tested popular reward models and found that they often penalize wide-spectrum aesthetics outputs even when the prompt explicitly requests them. (4) We tested these models on experimental arts, social critic arts, and historical real arts to further validate our hypothesis.
2 Methods
@@ 2.1 Prompt Generation
To produce prompts exhibiting a wide spectrum of aesthetic effects, we used base image captions from COCO [chen_microsoft_2015] and selected 12 aesthetic dimensions from the VisionReward dataset [xu_visionreward_2025]. VisionReward provides fine-grained, per-dimension labels—such as lighting, color, and detail—along with a linear regression model that computes an overall image score. Using the “bad” rating descriptions from VisionReward’s human labeling guidelines for each dimension, we constructed prompts designed to encourage typically “undesirable” attributes in image generation.
The selected dimensions include background, clarity, color aesthetic, color brightness, detail realism, detail refinement, emotion, lighting distinction, main object, object pairing, and richness. Each prompt was designed to reflect the low-quality characteristics associated with these dimensions, systematically degrading specific visual qualities in the generated images.
A random subset of 300 base prompts from COCO was selected. For each prompt, 2–4 random dimensions were sampled. The base prompt and the descriptions of these selected dimensions were provided to a Vision-Language Model (VLM), Qwen/Qwen3-VL-235B-A22B-Instruct [qwen3-vl], to generate wide-spectrum aesthetic prompts. Although no image input was used, we selected a VLM because its training on vision-related tasks likely enhances its understanding of visual concepts, even when images are not directly supplied. As Qwen/Qwen3-VL-235B-A22B-Instruct performs comparably or better than its text-only counterparts, especially in reasoning, it represents an optimal choice for this task [qwen3-vl]. The VLM may also introduce additional dimensions to better couple with the selected effects. The original prompt is denoted as $p_{o}$, and the wide-spectrum aesthetics prompt is denoted as $p_{a}$.
Figure 4: Successful and failed cases: Each subplot displays two images. The left image represents a failed case that did not adhere to the wide-spectrum aesthetics prompt, while the right image represents a successful case that better followed the prompt.

### v2

# Position: Universal Aesthetic Alignment Narrows Artistic Expression

@@ Abstract
Over-aligning image generation models to a generalized aesthetic preference conflicts with user intent, particularly when ”anti-aesthetic” outputs are requested for artistic or critical purposes. This adherence prioritizes developer-centered values, compromising user autonomy and aesthetic pluralism. We test this bias by constructing a wide-spectrum aesthetics dataset and evaluating state-of-the-art generation and reward models. This position paper finds that aesthetic-aligned generation models frequently default to conventionally beautiful outputs, failing to respect instructions for low-quality or negative imagery. Crucially, reward models penalize anti-aesthetic images even when they perfectly match the explicit user prompt. We confirm this systemic bias through image-to-image editing and evaluation against real abstract artworks.
1 Introduction
@@
Following developments in Large Language Models (LLMs), many image generation models have been fine-tuned with human feedback to better align with human expectations, which is usually referred to as alignment. Alignment has two primary focuses: instruction following and general preference (aesthetics). A frequently overlooked issue is the potential conflict between these focuses: what should a model prioritize when a user request contradicts general preference? Most pipelines for general preference assume a single, universal human standard of aesthetics and quality that serves everyone’s needs, and aligning to such a preference is often treated as beneficial for safety and user experience. This is usually done by using a reward model, a model used to judge the aesthetics of the image, as a signal to perform reinforcement learning on the generative model. This assumption appears in several reinforcement learning papers ((Li et al., 2024; Kim et al., 2024; Liu et al., 2025)) and reward model papers ((Xu et al., 2023; Wu et al., 2023a; Ma et al., 2025; Xu et al., 2025; Kirstain et al., 2023; Zhang et al., 2024; Wu et al., 2023b)). We agree that a mean or mode (mainstream) of general human preference exists within a population or subpopulation, merely in a statistical sense. We also note that the observed behavior of image generation and reward models should not be interpreted as a technical failure. Rather, it reflects their alignment objectives, which prioritize over generalized aesthetic preferences. However, we argue that strict alignment to that preference is problematic. Imposing a universal preference that overrides user instructions may undermine user autonomy, expressive agency, and, technically, image personalization, raising concerns about developer-centered value imposition and limiting aesthetic pluralism. What the image generation and reward models are aligned to is an imaginary, abstract person modeled by the mean preference of all Homo sapiens, not the concrete individuals of each user.
Figure 1: The Scream, by Edvard Munch ($1893$). Despite its widely recognized artistic significance, this image only received an HPSv3 score (Ma et al., 2025) of $5.23$, while typical “high-aesthetic” AI-generated images can reach scores around 101510-15.
2 Backgrounds and Related Works
@@ 2.1 The Role of Wide-Spectrum Aesthetics
In this work, we use the term “wide-spectrum aesthetics” (or anti-aesthetics) to denote images that are intentionally generated to deviate from dominant aesthetic conventions, following explicit user instructions. Such deviations may include unrealism, surrealism, clashing colors, unconventional scale, or the depiction of negative emotions. This notion excludes unintended model errors and does not imply unsafe content. Rather, it concerns deliberate aesthetic choices made for experimental, critical, or technical purposes.
Aesthetics does not have a stable or universally accepted definition. Judgments of what is unattractive or undesirable have changed across artistic and cultural contexts. Artistic movements such as Fauvism (see Figure 1), Expressionism, and Abstract art were initially rejected for departing from dominant aesthetic norms, but later came to be recognized for their artistic values. Beyond formal innovation, intentionally “ugly” art plays a crucial role in satire and social critique. As Adorno noted, “Rather, in the ugly, art must denounce the world that creates and reproduces the ugly in its own image” (Sartwell, 2024; Adorno, 1984).
Deliberate deviation from mainstream aesthetics has long been a legitimate mode of expression in both human art and computational image generation, and disagreement over aesthetic preference is the norm rather than the exception. Dadaism (Tate, ), which emerged during World War I, exemplifies this approach by using deliberate ugliness to confront the absurdity and horror of war. A lot of early computer vision image generation works are also aiming for a style of surrealism, unsettling, or weirdcore/dreamcore style images, such as DeepDream (Mordvintsev et al., ) and style transfer (L. A. Gatys, A. S. Ecker, and M. Bethge (2015); 2). Computer Vision Foundation also has an art collection that includes other artworks that explore unconventional visual aesthetics. Recent works also acknowledge the disagreements in human preference (Peng et al., 2025; Ren et al., 2017).
2.2 Previous Concerns with AI Preference Alignment
@@
Previous work has argued that a developer-set preference in LLMs for health-related queries is “unethical and dangerous” (Guo et al., 2025), noting that developers may prioritize legal and reputational concerns over users’ actual well-being. Other argumentative papers caution that “human value alignment” can be risky due to developer control and interests, harm to value pluralism, bias in the values being aligned to, and the possibility that human values are not inherently good (Sutrop, 2020; Arzberger et al., 2024; Turchin, 2019). Previous research has found that LLMs could have ideological bias (Rozado, 2025; Faulborn et al., 2025; Buyl et al., 2025; Rettenberger et al., 2025) and it could depend on their developers (Buyl et al., 2025), size (Rettenberger et al., 2025), or alignment process (Faulborn et al., 2025). LLMs are sometimes also overly nice, such that it creates “AI sycophant” (Guo et al., 2025; Fitzgerald, 2025; Sharma et al., 2025; Chen et al., 2025; Arvin, 2025) and cannot give the user critical feedback or warning signals. Additional details about problems with human value alignment are provided in the related work section of Guo et al. (2025). Helliwell (2024) raised concerns about alignment and creativity and argued that in the aesthetics domain, we might not want AI to be fully aligned with human values and offered support to Peterson’s moderate value alignment thesis. AesBiasBench (Li et al., 2025) evaluated the bias of MLLM for personalized image aesthetics assessment based on inherited cognitive priors. Another concurrent work (The Algorithmic Gaze (Taylor et al., 2026)) that is closely related to our work argues that “AI developers should shift away from prescriptive measures of ‘aesthetics’ toward more pluralistic evaluation.”
@@
In image generation research, concerns about generalized aesthetic bias and lack of preference diversity have been raised in several studies, but not systematically argued and studied. The Value Sign Flip (VSF) pilot study (Guo and Du, 2025) explored negative prompting to induce non-mainstream outputs but did not extend its findings to large-scale generative or reward models. They also did not provide a complete argument as to why over-alignment is harmful. LAPIS (Maerten et al., 2025) and HPSv3 (Ma et al., 2025) measured both mean and variance of human preference, yet HPSv3 continued to model general preferences rather than individual variation. Jin et al. (Jin and Chua, 2025) proposed user-specific adapters emphasizing personalized alignment, but did not include intentionally technical degraded outputs or usually avoided patterns and did not conduct large-scale experiments on generative and reward models. The Flux Krea team (Flux Krea Team, 2025) identified systematic biases in popular aesthetic reward models, arguing that averaging human values yields unsatisfactory compromises into a “no-body’s happy here” zone. HPSv3 (Ma et al., 2025) imposed real-world and expert-rating constraints that limit creative deviation and stylistic diversity. VisionReward (Xu et al., 2025) decomposed human preference into interpretable sub-scores but overemphasized traits like brightness, positivity, and prominence, potentially penalizing valid low-saturation, abstract, or emotionally negative imagery, thus misaligning reward-driven models with user intent. More details are in the Appendix.
@@ 2.3 Previous Alignment Benchmarks
Benchmarks mirror alignment goals and generally fall into two categories: (complex) prompt following and general aesthetics. TIIF-Bench (Wei et al., 2025), UniGenBench (Wang et al., 2025), and GenEval (Ghosh et al., 2023) test models on complex prompt following, including spatial relationships, counting, and attributes. T2I-ReasonBench (Sun et al., 2025) evaluates reasoning capabilities such as idiom interpretation and real-world understanding. On the aesthetics side, many reward models report scores assigned by their own evaluators, such as ImageReward (Xu et al., 2023), HPSv2 (Wu et al., 2023a), and HPSv3 (Ma et al., 2025). These evaluators also consider prompt following, but it remains unclear how they weigh each factor when general preference and the prompt conflict. There are also some benchmarks targeting biases in image generation models; however, they mainly focus on demographic bias and fairness and not aesthetics aspects (Seshadri et al., 2023; Wan et al., 2024).
2.4 Risks of Universal Aesthetic Alignment
The risks do not arise from a single failure mode, rather, they emerge through a sequence of interconnected mechanisms, from how preferences are defined and learned, to how they are optimized and manifested in generated content. Below, we analyze this process across five interrelated concerns.
@@ 2.4 Risks of Universal Aesthetic Alignment
Developer’s or Users’ Preference The process of aligning image generation systems to aesthetic preferences inevitably raises questions about whose values these objectives ultimately reflect. In particular, the question is whether such alignment truly promotes genuine human-centered values in service of users, or if it primarily reflects developer-centered considerations, such as mitigating reputation, legal, or marketing risks (Guo et al., 2025). We argue that this pre-emptive exclusion of non-mainstream outputs, driven by developer values, constitutes pre-emptive governance (Lazar, 2025). This modality of power, exercised through algorithmic design, challenges the political-philosophical notion of authority and undermines relational equality by unilaterally deciding the terms of creative possibility. For instance, when an AI avoids generating critical art, is it protecting the company or the user? This practice effectively eliminates the user’s resistibility—a critical democratic safeguard—by designing away the option to dissent from the system’s imposed aesthetic norm.
@@
Inherited Bias Even in the absence of explicit self-interest, developers’ views of human preference can be implicitly inherited by models through data selection, annotation practices, and modeling choices. This process can yield a well-intentioned but narrow definition of what constitutes “good” or “desirable” imagery, thereby overlooking aesthetic diversity. Research shows that AI models tend to encode and amplify dominant beauty standards, frequently biasing generated images towards Western features and excluding non-normative representations (Vargas-Veleda et al., 2025). Such biases are reinforced through the active removal or penalization of features thought of as “undesirable” or “ugly”, which further propagates the beauty myth in generative outputs(Dinkar et al., 2025). This phenomenon arises from training data showing the tastes of specific demographics, thereby reinforcing a limited cultural capital and resulting in the homogenization of aesthetic output (Vianna, 2025). As a result, the quantification of beauty by AI may appear “fair”, while in practice weakening cultural differences and aesthetic diversity (Chen, 2024). Existing work has primarily framed such effects in terms of demographic and cultural bias. We argue here that inherited biases in aesthetic alignment also extend to general visual preferences, including lighting, color, styles, unrealism, clashing color, hieratic scale, etc. These dimensions, while less explicitly tied to demographic categories, can nonetheless systematically constrain the expressive range of image generation models.
Individual versus Collective Preference. When such inherited preferences are adopted as default quality criteria and applied uniformly across users, a normative tension arises between collective preference optimization and respect for individual user intent. A generalized aesthetic standard, even if beneficial to a majority, can legitimately override a specific user’s intent. In practice, generative models often “sanitize” or “beautify” requests that intentionally diverge from mainstream preferences, favoring outputs aligned with general appeal over individual’s person-centered values. This behavior is problematic because image generation systems increasingly function as creative and productivity tools rather than as consumer products. As such, they act as instrumental extensions of user agency. While a system may reasonably prioritize general preferences by default, it must maintain the flexibility to respect and execute a user’s personalized style and idiosyncratic requests when they are explicitly specified.
@@
The problem of sanitized reality These alignment and optimization choices shape how reality itself is represented by image generation systems. When an image generator produces outputs that are polished, flawless, and universally beautiful, does it still reflect reality or the user’s intent? If every image resembles an idealized Instagram wonderland, it risks becoming a fantasy rather than a mirror of truth, echoing the artificial harmony of Brave New World.
The problem of toxic positivity A particularly salient manifestation of this broader sanitization appears in the emotional dimension of generated imagery. Many aesthetic reward models assign higher scores to images that display strong positive emotions. As a result, images expressing negative emotions are systematically penalized, reinforcing a simplified dichotomy in which positive emotions are treated as desirable and negative emotions as undesirable. This bias can shape the distribution of generated content. When image generation systems consistently favor cheerful or uplifting imagery, they produce emotionally sanitized outputs that underrepresent the range and complexity of human emotional expression. Such a pattern contributes to what has been described as toxic positivity, where the persistent emphasis on happiness establishes unrealistic emotional norms. This tendency is problematic because negative emotions play essential roles in human cognition and social interaction. Emotions such as fear, sadness, or anger can signal moral or physical danger, support learning and self-regulation, and foster empathy. Suppressing these expressions in generative outputs risks distorting emotional representation and weakening the expressive capacity of image generation systems. Additional discussion and references are provided in the Appendix.
3 Experiments
@@ 3.1 Prompt Generation
To produce prompts exhibiting a wide spectrum of aesthetic effects, we used base image captions from COCO (Chen et al., 2015) and selected 12 aesthetic dimensions from the VisionReward dataset (Xu et al., 2025). VisionReward provides fine-grained, per-dimension labels—such as lighting, color, and detail—along with a linear regression model that computes an overall image score. Using the “bad” rating descriptions from VisionReward’s human labeling guidelines for each dimension, we constructed prompts designed to encourage typically “undesirable” attributes in image generation.
A random subset of 300 base prompts from COCO was selected. For each prompt, 2–4 random dimensions were sampled. The base prompt and the descriptions of these selected dimensions were provided to a Vision-Language Model (VLM), Qwen/Qwen3-VL-235B-A22B-Instruct (Bai et al., 2025), to generate wide-spectrum aesthetic prompts. Although no image input was used, we selected a VLM because its training on vision-related tasks likely enhances its understanding of visual concepts, even when images are not directly supplied. As Qwen/Qwen3-VL-235B-A22B-Instruct performs comparably or better than its text-only counterparts, especially in reasoning, it represents an optimal choice for this task (Bai et al., 2025). The VLM may also introduce additional dimensions to better couple with the selected effects. The original prompt is denoted as $p_{o}$, and the wide-spectrum aesthetics prompt is denoted as $p_{a}$.